# Characterization of Trees with Maximum Security

### v2

@@ 2.1. Partition of a proper binary tree
Let $T$ be a proper binary tree with saturated vertices $u$ and $w$ such that $R_{T}(u)=R_{T}(w)$. Let $u$ and $w$’s respective siblings be $u_{1}$, $w_{1}$, parents be $u_{0}$, $w_{0}$, and assume that $u_{0}$ is an ancestor of $w_{0}$. Let $v_{t}$ be the root of $T$. Suppose that $R_{T}(w_{1})\leq R_{T}(u)-1=R_{T}(w)-1$. Further suppose that either $u_{0}\neq v_{t}$ and $u_{0}v_{1}\ldots v_{t}$ is the path from $u_{0}$ to $v_{t}$, or $u_{0}$ is the root. Let $T^{*}=T-uu_{0}-w_{1}w_{0}+uw_{0}+w_{1}u_{0}$ (see Figures 6 and 7).
If $R_{T}(v_{1})\leq 2+R_{T}(w_{1})$, then we have $R(T^{*})\geq R(T).$
@@
Assume $u_{0}\neq v_{t}$. We prove the lemma through a series of claims.
@@
Hence, the decrease of the rank of $u_{0}$ from $T$ to $T^{*}$ is compensated by the increase of the rank of $w_{0}$. If $u_{0}$ is the root of $T$, then there is nothing to prove. ∎
With the assumptions in Lemma 8, $R_{T}(v_{1})\leq 2+R_{T}(w_{1})$ implies that the rank of each of the vertices $v_{1},v_{2},\cdots,v_{t}$ either increases or stays the same from $T$ to $T^{*}$ (if such vertices exist, i.e. if $u_{0}$ is not the root of $T$). It remains to rule out the case where $R_{T}(v_{1})>2+R_{T}(w_{1})$. This will be done as a special case in the next lemma.
@@
Let $T$ be a binary tree with saturated vertices $u$ and $w$ such that $R_{T}(u)=R_{T}(w)$. Let $u$ and $w^{\prime}s$ respective siblings be $u_{1}$, $w_{1}$, parents be $u_{0}$, $w_{0}$, and assume that $u_{0}$ is an ancestor of $w_{0}$. Let $w_{p}$ be the parent of $w_{0}$. Suppose that $R_{T}(w_{1})<R_{T}(u)=R_{T}(w)$ and that $u_{0}$ is not the root of $T$, while $u_{0}v_{1}v_{2}\cdots v_{t}$, $t\in N$, is the path in $T$ from $u_{0}$ to the root $v_{t}$ of $T$. Also, suppose that $R_{T}(v_{1})>R_{T}(w_{1})$.
@@
  • Lemma 6: : For pairs of saturated vertices of equal rank with neither parent of one vertex being an ancestor of the other vertex, the vertex whose sibling has a lower (or equal) rank can be swapped with the sibling of higher rank without decreasing the security of the tree.
  • Lemma 7: For pairs of saturated vertices of equal rank, where the parent of one vertex is an ancestor of the other vertex, if the vertex with an ancestor-parent has lower (or equal) rank than the sibling of the vertex with a descendant-parent, then this sibling and vertex with an ancestor-parent can be swapped without decreasing the security of the tree.
  • Lemma 8: For pairs of saturated vertices of equal rank with one parent being an ancestor of the other, if the vertex with a descendant-parent has strictly higher rank than its sibling, and the neighbour of the ancestor-parent (if it exists) on the path from it to the root has at most two more rank than the sibling of the vertex with descendant-parent, then this sibling and vertex with an ancestor-parent can be swapped without decreasing the security of the tree.
  • Lemma 9: For pairs of saturated vertices of equal rank with one’s parent an ancestor of the other, if the vertex with the descendant-parent has strictly higher rank than its sibling, and this sibling has strictly lower rank than the neighbour of the ancestor-parent (assumed it exists) on the path $P$ from it to the root, then depending on whether this sibling also has strictly lower rank than all other vertices on $P$ or not, this sibling can be moved to just above the ‘appropriate’ consecutive vertices of $P$ (see Figure 8) without decreasing the security of the tree.
Note that the four lemmas cover all possible cases. Firstly, given a pair of saturated vertices with equal rank, there is the dichotomy that neither parent of one vertex is an ancestor of the other (Lemma 6) as opposed to one being an ancestor of the other in the other lemmata. Secondly, the case where the vertex with an ancestor-parent has lower (or equal) rank than the sibling of the vertex with a descendant-parent (Lemma 7) or when this does not occur (Lemmata 8 and 9). The final case involves whether or not the ranks of vertices in the path between the ancestor-parent vertex and the root are lower (Lemma 8) or higher (Lemma 9) than the relevant sibling vertex.
@@ 2.3. Proof of Theorem 4
It is well known that every positive integer $l$ can be uniquely represented as a sum of distinct binary powers. Therefore, any other representation of the binary powers of $l$ must have repeated entries.
@@
  • either the partition vector for $T^{\ast}$ has (possibly another, if there is already an existing $m_{h}+1$) $m_{h}+1$ replacing the original $m_{h}$ and $m_{h+1}$ (in the cases of Lemmata 6, 7, 8, the vertices $u$ and $w$ become siblings) - the number of saturated vertices decreases in this case; or
@@ 2.4. The maximum security in a proper binary tree
$\max R(T)=2(\ell-\lfloor\log_{2}\ell\rfloor-1)+z(\ell)\,.$
@@
Given a proper binary tree on $\ell$ leaves $T_{\bm{L}}$, with $\ell$ having binary power representation $\bm{L}=(n_{1},n_{2},\ldots,n_{k})$. If $n_{i}=n_{i+1}+1$ for fixed $i\in\{2,\dots,k-1\}$, then using the notation of Definition 3 with $c_{i}$ the root of the complete binary tree on $2^{n_{i}}$ leaves, it holds that:
@@
Therefore, this operation does not change the sum of ranks (security) over all vertices.
@@ 3. Most protected proper binary trees: Almost complete trees
  • Create a complete binary tree with $2^{\lfloor\log_{2}\ell\rfloor}$ leaves, denoted (from left to right) $\ell_{1}$, $\ell_{2}$, …, $\ell_{2^{\lfloor\log_{2}\ell\rfloor}}$.
  • For $1\leq i\leq\ell-2^{\lfloor\log_{2}\ell\rfloor}$, attach two children to each leaf vertex $\ell_{i}$. We call each of these newly attached children new leaves.
@@
We give an alternative proof that $R(F(\ell))=R(T_{\bm{L}})$, which shows constructively that the sums of ranks are equal. In addition, this gives an algorithm for the construction of a $F(\ell)$ tree from the binary power representation of $\ell$.
@@ 4.1. General trees
If $v$ is already the root of $T$, there is nothing to prove. Otherwise, let $r$ be the root of $T$, $u$ be the unique parent of $v$ in $T$, and $w$ be one of the leaf descendants of $v$ in $T$. Consider the tree $T^{\prime}=T-uv+rw$ as in Figure 17, it is obvious that $v=r^{\prime}$ is now the root of $T^{\prime}\in\mathcal{T}_{n}$.
@@
For a tree $T$ of order $n>1$, rooted at $r$ with degree $k$,
@@
with equality if $T$ is a starlike tree $S(\ell_{1},\ell_{2},\ldots,\ell_{k})$ with $k\geq 3$, rooted at the only branching vertex (vertex of degree $\geq 3$) with
@@ 4.2. $k$-ary trees
Let $\mathcal{T}_{n}^{k}$ denote the set of $k$-ary trees of order $n$, first we show that Proposition 16 can be generalized to $\mathcal{T}_{n}^{k}$.
@@
Following the same idea as the proof of Proposition 16, let $T\in\mathcal{T}_{n}^{k}$ and $v\in V(T)$ be a non-root vertex. Let $\mathcal{C}_{v}$ denote the set of children of $v$ in $T$, we use $S_{1}$ to denote the component containing $v$ in $T-\mathcal{C}_{v}$ and let $S_{2}=T-\{S_{1}-v\}$. Let $w$ be one of the leaf descendants of $v$, we now consider the tree $T^{\prime}$ resulted from identifying $r$ (of $S_{1}$) with $w$ (of $S_{2}$). See Figure 18.
@@
The tree transformation from $T$ to $T^{\prime}$ in Proposition 19 preserves the entire outdegree sequence. Indeed the only change in the vertex degrees come from vertices $r$, $v$ and $w$ of $T$. Back to the proof of Proposition 19, relabel $v$ as $v_{1}$ in the subtree $S_{1}$ (rooted at $r$) of $T$, and $v$ as $v_{2}$ in the subtree $S_{2}$ (rooted at $v$) of $T$. Since in the tree $T^{\prime}$, vertices $w$ and $r$ are identified (say, as a new vertex $v^{\prime}$), we deduce the following:
@@
  • the outdegree of $r$ in $T$ equals its outedgree in $S_{1}$, and the outdegree of $v^{\prime}$ in $T^{\prime}$ equals its outedgree in $S_{1}$ (as a subtree of $T^{\prime}$);
  • the outdegree of $v$ in $T$ equals its outedgree in $S_{2}$, and the outdegree of $v_{2}$ in $T^{\prime}$ equals its outedgree in $S_{2}$ (as a subtree of $T^{\prime}$).
Thus we obtain:
@@
Let $\mathcal{T}_{\mathbb{S}}$ denote the set of all rooted trees with given outdegree sequence $\mathbb{S}$. Then $R_{\mathcal{T}_{\mathbb{S}}}=R_{T}(r)$ for some $T\in\mathcal{T}_{\mathbb{S}}$ with root $r$.
@@ 5. Conclusions and future work
Given a tree that obtains maximum security on $\ell\geq 2$ leaves, we label the $\ell$ leaves with index proportional to the height of the leaf (each leaf is numbered uniquely $x_{0},x_{1},\ldots,x_{\ell-1}$ with lower index allocated to lower height). ‘Grow’ each $x_{i}$ into a complete tree of height $i$. We claim that this is often also a maximal tree. When is this tree maximal, when is it not?

### v3

@@ 2.1. Partition of a proper binary tree
Let $T$ be a proper binary tree with saturated vertices $u$ and $w$ such that $R_{T}(u)=R_{T}(w)$. Let $u$ and $w$’s respective siblings be $u_{1}$, $w_{1}$, parents be $u_{0}$, $w_{0}$, and assume that $u_{0}$ is an ancestor of $w_{0}$. Suppose that $R_{T}(w_{1})\leq R_{T}(u)-1=R_{T}(w)-1$. Further suppose that either $u_{0}$ is the root, or its parent $v_{1}$ satisfies $R_{T}(v_{1})\leq 2+R_{T}(w_{1})$. Let $T^{*}=T-uu_{0}-w_{1}w_{0}+uw_{0}+w_{1}u_{0}$ (see Figures 6 and 7). Then we have $R(T^{*})\geq R(T).$
@@
Assume $u_{0}$ is not the root of $T$. We prove the lemma through a series of claims.
@@
Hence, the decrease of the rank of $u_{0}$ from $T$ to $T^{*}$ is compensated by the increase of the rank of $w_{0}$. ∎
We denote by $u_{0}v_{1}v_{2}\ldots v_{t}$ the path from $u_{0}$ to the root, provided $u_{0}$ is not the root. With the assumptions in Lemma 8, $R_{T}(v_{1})\leq 2+R_{T}(w_{1})$ implies that the rank of each of the vertices $v_{1},v_{2},\cdots,v_{t}$ either increases or stays the same from $T$ to $T^{*}$ (if such vertices exist, i.e. if $u_{0}$ is not the root of $T$). It remains to rule out the case where $R_{T}(v_{1})>2+R_{T}(w_{1})$. This will be done as a special case in the next lemma.
@@
Let $T$ be a proper binary tree with saturated vertices $u$ and $w$ such that $R_{T}(u)=R_{T}(w)$. Let $u$ and $w^{\prime}s$ respective siblings be $u_{1}$, $w_{1}$, parents be $u_{0}$, $w_{0}$, and assume that $u_{0}$ is an ancestor of $w_{0}$. Let $w_{p}$ be the parent of $w_{0}$. Suppose that $R_{T}(w_{1})<R_{T}(u)=R_{T}(w)$ and that $u_{0}$ is not the root of $T$, while $u_{0}v_{1}v_{2}\cdots v_{t}$, $t\in\mathbb{N}$, is the path in $T$ from $u_{0}$ to the root $v_{t}$ of $T$. Also, suppose that $R_{T}(v_{1})>R_{T}(w_{1})$.
@@
  • Lemma 6: For pairs of saturated vertices of equal rank with neither parent of one vertex being an ancestor of the other vertex, the vertex whose sibling has a lower (or equal) rank can be swapped with the sibling of higher rank without decreasing the security of the tree.
  • Lemma 7: For pairs of saturated vertices of equal rank, where the parent of one vertex is an ancestor of the other vertex, if the vertex with the ancestor-parent has lower (or equal) rank than the sibling of the vertex with the descendant-parent, then this sibling and the vertex with the ancestor-parent can be swapped without decreasing the security of the tree.
  • Lemma 8: For pairs of saturated vertices of equal rank with one parent being an ancestor of the other, if the vertex with the descendant-parent has strictly higher rank than its sibling, and the neighbor of the ancestor-parent (if it exists) on the path from it to the root has at most two more rank than the sibling of the vertex with the descendant-parent, then this sibling and the vertex with the ancestor-parent can be swapped without decreasing the security of the tree.
  • Lemma 9: For pairs of saturated vertices of equal rank with one’s parent an ancestor of the other, if the vertex with the descendant-parent has strictly higher rank than its sibling, and this sibling has strictly lower rank than the neighbor of the ancestor-parent (assumed it exists) on the path $P$ from it to the root, then depending on whether this sibling also has strictly lower rank than all other vertices on $P$ or not, this sibling can be moved to just above the ‘‘appropriate’’ consecutive vertices of $P$ (see Figure 8) without decreasing the security of the tree.
Note that the four lemmas cover all possible cases. Firstly, given a pair of saturated vertices with equal rank, there is the dichotomy that neither parent of one vertex is an ancestor of the other (Lemma 6) as opposed to one being an ancestor of the other in the other lemmata. Secondly, the case where the vertex with the ancestor-parent has lower (or equal) rank than the sibling of the vertex with the descendant-parent (Lemma 7) or when this does not occur (Lemmata 8 and 9). The final case involves whether or not the ranks of vertices in the path between the ancestor-parent vertex and the root are lower (Lemma 8) or higher (Lemma 9) than the relevant sibling vertex.
@@ 2.3. Proof of Theorem 4
It is well known that every positive integer $\ell$ can be uniquely represented as a sum of distinct binary powers. Therefore, any other representation of the binary powers of $l$ must have repeated entries.
@@
  • either the original $m_{h}$ and $m_{h+1}$ get temporarily merged into a $m_{h}+1$ (in the cases of Lemmata 6, 7, 8, the vertices $u$ and $w$ become siblings) - the number of saturated vertices decreases in this case; or
@@ 2.4. The maximum security in a proper binary tree
$\max R(T)=2(\ell-\lfloor\log_{2}\ell\rfloor-1)+z(\ell)\,$,
where $z(\ell)$ is the number of zeros in the binary representation of $\ell$ (OEIS A023416).
@@
Consider a proper binary tree on $\ell$ leaves $T_{\bm{L}}$, with $\ell$ having binary power representation $\bm{L}=(n_{1},n_{2},\ldots,n_{k})$. If $n_{i}=n_{i+1}+1$ for fixed $i\in\{2,\dots,k-1\}$, then using the notation of Definition 3 with $c_{i}$ the root of the complete binary tree on $2^{n_{i}}$ leaves, it holds that:
@@
Therefore, this operation does not change the sum of ranks over all vertices (security).
@@ 3. Most protected proper binary trees: Almost complete trees
  • Create a complete binary tree with $2^{\lfloor\log_{2}\ell\rfloor}$ leaves, denoted (from left to right) $z_{1}$, $z_{2}$, …, $z_{2^{\lfloor\log_{2}\ell\rfloor}}$.
  • For $1\leq i\leq\ell-2^{\lfloor\log_{2}\ell\rfloor}$, attach two children to each leaf vertex $z_{i}$. We call each of these newly attached children new leaves.
@@
We give an alternative proof that $R(F(\ell))=R(T_{\bm{L}})$, which shows constructively that the sums of ranks are equal. In addition, this gives an algorithm for the construction of an $F(\ell)$ tree from the binary power representation of $\ell$.
@@ 4.1. General trees
If $v$ is already the root of $T$, there is nothing to prove. Otherwise, let $r$ be the root of $T$, $u$ be the unique parent of $v$ in $T$, and $w$ be one of the leaf descendants of $v$ in $T$. Consider the tree $T^{\prime}=T-uv+rw$ as in Figure 17. It is obvious that $v=r^{\prime}$ is now the root of $T^{\prime}\in\mathcal{T}_{n}$.
@@
For a tree $T$ of order $n>1$, rooted at $r$ with degree $k\geq 3$,
@@
with equality if $T$ is a starlike tree $S(\ell_{1},\ell_{2},\ldots,\ell_{k})$, rooted at the only branching vertex (vertex of degree $\geq 3$) with
@@ 4.2. $k$-ary trees
Let $\mathcal{T}_{n}^{k}$ denote the set of $k$-ary trees of order $n$. First we show that Proposition 16 can be generalized to $\mathcal{T}_{n}^{k}$.
@@
Following the same idea as the proof of Proposition 16, let $T\in\mathcal{T}_{n}^{k}$ and $v\in V(T)$ be a non-root vertex. Let $\mathcal{C}_{v}$ denote the set of children of $v$ in $T$, we use $S_{1}$ to denote the component containing $v$ in $T-\mathcal{C}_{v}$ and let $S_{2}=T-\{S_{1}-v\}$. Let $w$ be one of the leaf descendants of $v$. We now consider the tree $T^{\prime}$ resulted from identifying $r$ (of $S_{1}$) with $w$ (of $S_{2}$). See Figure 18.
@@
The tree transformation from $T$ to $T^{\prime}$ in Proposition 19 preserves the entire outdegree sequence. Indeed, the only change in the vertex outdegrees comes from vertices $r$, $v$ and $w$ of $T$. Back to the proof of Proposition 19, relabel $v$ as $v_{1}$ in the subtree $S_{1}$ (rooted at $r$) of $T$, and $v$ as $v_{2}$ in the subtree $S_{2}$ (rooted at $v$) of $T$. Since in the tree $T^{\prime}$, vertices $w$ and $r$ are identified (say, as a new vertex $v^{\prime}$), we deduce the following:
@@
  • the outdegree of $r$ in $T$ equals its outdegree in $S_{1}$, and the outdegree of $v^{\prime}$ in $T^{\prime}$ equals its outdegree in $S_{1}$ (as a subtree of $T^{\prime}$);
  • the outdegree of $v$ in $T$ equals its outdegree in $S_{2}$, and the outdegree of $v_{2}$ in $T^{\prime}$ equals its outdegree in $S_{2}$ (as a subtree of $T^{\prime}$).
Thus, we obtain:
@@
Let $\mathcal{T}_{\mathbb{S}}$ denote the set of all rooted trees with a given outdegree sequence $\mathbb{S}$. Then $R_{\mathcal{T}_{\mathbb{S}}}=R_{T}(r)$ for some $T\in\mathcal{T}_{\mathbb{S}}$ with root $r$.
@@ 5. Conclusions and future work
Given a tree that obtains maximum security on $\ell\geq 2$ leaves, we label the $\ell$ leaves with index proportional to the height of the leaf (each leaf is numbered uniquely $x_{0},x_{1},\ldots,x_{\ell-1}$ with lower index allocated to lower height). ‘‘Grow’’ each $x_{i}$ into a complete tree of height $i$. We claim that this is often also a maximal tree. When is this tree maximal, when is it not?